North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog 8th and 9th December 2020

8th and 9th December 2020

On a cold, misty, drab, Tuesday morning at Grosmont, Bill Dobson, Nigel Hall, Nigel Bill and Ian Pearson arrived to find smoke issuing from the Q6 chimney. A warming fire was lit on Monday afternoon then relit on Tuesday morning (see photo).

63395 boiler out of frames steam test on 8 December 2020 -  Nigel Hall

Away from that excitement, Nigel B and Ian continued to clean between the frames. Nigel B also dressed some flaking paint from the weigh shaft area and applied a coat of red oxide on the bare metal. The lifting jacks had arrived back from Pickering on Monday, but, when tested, they would not work. An electrical fault was found in the control unit which Chris White (electrician) rectified during the morning. So, before lunch, all four jacks were moved into position by Nigel B and Ian, Chris Cubitt, Sean Bowler, Andrew Jeffery and Dave Cholmondeley. Moving these jacks is very heavy work, but they were now set ready for Sean to do some final adjustments and measurements to the horn guides the next day. Bill started rubbing down the vacuum pipes which Ed cleaned last week, while Nigel H gave the back plate, cab steps and tea tray a coat of black gloss paint. Later Nigel H dressed the steam heat pies with a flapper wheel and gave them a coat of black primer as well as giving the vacuum pipes a coat of red oxide where required.

After lunch the steam test on the boiler had to be abandoned for various reasons, and which were to be sorted out the next day. These included the absence of blanks on the header: these were found upstairs in the Deviation Shed workshop. Ian made some new gaskets for them and they were to be fitted tomorrow. Also the packing under the foundation ring area required some alteration, and the blower pipe from the manifold to the operating valve required fitting. Bill started this job just before home time and was to continue the next day. It now looks as if the boiler inspector will be unable to come for the formal out of frames steam test until January 2021.

The weather on Wednesday was more of the same, although it brightened up later for Steve Hyman, Nigel Hall, Bill Dobson, Ian Pearson and, at lunchtime, Jonathan Bradley.

Steve fitted the blower supply pipe from the manifold to the blower valve on the back head and Ian fitted the two blanks to the header in the smokebox. One stud was missing so Bill asked Barry Nesom to make a new one. This was fitted and both blanks are now in situ on the header. Ian removed the gauge frame protectors and these are now stored in a black plastic bucket in the large locker upstairs in Deviation shed. They require some attention - new springs,cleaning and painting. Nigel H continued to paint - another coat of black gloss on the steam heat pipes, red gloss on the vacuum pipes, and eventually got down to the frames later in the afternoon. He then painted all the parts that had been painted in red oxide yesterday with red gloss.

After lunch Steve and Jon worked outside on back head parts, dressing by removing old flaking paint but there is still quite an amount to do so Nigel H is going to acquire some dressing disks.

Bill went to local steel stockists to buy a steel beam of 6 feet in length for fitting underneath the foundation ring as the wooden beam started to burn yesterday. Bill is arranging delivery to Grosmont sometime shortly. The water in the boiler is to be drained on Friday, then it will be jacked up, and the wooden beam replaced with the steel beam. Sean and Dave fitted the intermediate and leading wheel sets with new under keep pads. The four standing jacks were removed. Some final measurement checks on the horn guides showed all to be OK. The wheels are to be fitted back into the frames today (Thursday).

Locomotives in steam on both days were Lucie on steam heat duties and S15 on the Moorlander.